The following frequency table summarizes 50 data values. What is the 50th percentile of the data? Frequency 5 Value 3 3 5 7 8 9 10 2 12 13 14 15 16 3 Select the correct answer below:
15 16 Select the correct answer below: 7 2 Previous Page Next Page Page
Question 15 (1 point) If A and B are independent events, PIA)-0.13, and P(B)-0.85, what is P(AIB)? Select the correct answer below 0.11 0.00 o.85 ·0.98 0.72 0.13
Answer #1

1.) 50 th percentile is the value corresponding to cumulative frequency just greater than 50N/100 where N is the total frequency which is given to be 50.

So 50x50/100= 25.

Value corresponding to cumulative frequency just greater than 25 is 50th percentile of the data.

Value. Freq. Cumulative freq

1 . 5 5

2 2 7

3 4 11

4 3. 14

5. 2. 16

6. 4. 20

7. 6 . 26

8. 2. 28

9. 2. 30

10. 4. 34

11 . 1 35

12. 6. 41

13 . 2 . 43

14. 3 . 46

15 . 1. 47

16. 3. 50

Cumulative frequency just greater than 25 is 26 and value correspomding to it is 7 , so 50th percentile is 7

Ans . 7

Ques 1.5 ) P(A and B)= P(A) x P(B) = 0.13x0.85 = 0.1105 ( since A and B are independent)

P(A IB)= P( A and B)/ P(B) = 0.13

Ans. 0.13